# WeatherPrompt: Multi-modality Representation Learning for All-Weather Drone Visual Geo-Localization

### v2

@@ Abstract
Visual geo-localization for drones faces critical degradation under weather perturbations, e.g., rain and fog, where existing methods struggle with two inherent limitations: 1) Heavy reliance on limited weather categories that constrain generalization, and 2) Suboptimal disentanglement of entangled scene-weather features through pseudo weather categories. We present WeatherPrompt, a multi-modality learning paradigm that establishes weather-invariant representations through fusing the image embedding with the text context. Our framework introduces two key contributions: First, a Training-free Weather Reasoning mechanism that employs off-the-shelf large multi-modality models to synthesize multi-weather textual descriptions through human-like reasoning. It improves the scalability to unseen or complex weather, and could reflect different weather strength. Second, to better disentangle the scene and weather feature, we propose a multi-modality framework with the dynamic gating mechanism driven by the text embedding to adaptively reweight and fuse visual features across modalities. The framework is further optimized by the cross-modal objectives, including image-text contrastive learning and image-text matching, which maps the same scene with different weather conditions closer in the respresentation space. Extensive experiments validate that, under diverse weather conditions, our method achieves competitive recall rates compared to state-of-the-art drone geo-localization methods. Notably, it improves Recall@1 by +13.37% under night conditions and by 18.69% under fog and snow conditions.
@@ 1 Introduction
Drone visual geo-localization aims to match drone-view image with corresponding satellite views, supporting critical applications such as disaster response, urban surveillance, search-and-rescue, and environmental monitoring [1, 2, 3, 4, 5, 6]. However, variable weather conditions such as rain, fog and snow introduce noise, occlusions and low visibility, which severely distort image features [7, 8, 9] and cause conventional localization methods to suffer drastic performance degradation under extreme weather. Recent advances in cross-modal retrieval demonstrate that integrating natural language descriptions can substantially enhance the discrimination power of vision models, enabling better generalization in complex or ambiguous scenarios [10, 11, 12, 13]. Despite this progress, leveraging textual guidance for cross-weather drone geo-localization remains largely underexplored, especially considering the nuanced and dynamic nature of weather conditions encountered in the field. The ability of text to capture complex semantics and fine-grained details [14, 15] offers a promising avenue for cross-weather generalization. Addressing all-weather visual geo-localization task presents two primary challenges: (1) Limited Weather Labeling. Existing approaches typically perform domain-specific fine-tuning on a limited set of predefined weather labels (e.g., sunny, rainy). This closed-set paradigm fails to capture the continuous and combinatorial nature of real-world weather, thereby limiting model generalization to unseen or mixed conditions and preventing the exploitation of richer weather semantics. (2) Scene–Weather Feature Entanglement. Existing methods directly inject coarse pseudo-weather labels (e.g., rain, fog) into visual representations during training, leading to a severe entanglement between scene semantics and weather disturbances. Consequently, the model learns suboptimal representations under mixed or unseen weather conditions, fails to disentangle scene content from weather noise, and is severely limited in cross-weather generalization.
@@
Figure 1: Example of the proposed Chain-of-Thought description and matching. Our framework generates structured weather and spatial Text Description via stepwise reasoning. We leverage Off-the-shelf Visual Grounding Model (VGM), i.e., XVLM [16] to extract local region cues, which are integrated to further refine the matching process. Finally, we match images using weather description, global scene layout, and local region semantics to retrieve the corresponding satellite-view image.
@@
For the first limitation, we propose a training-free weather reasoning mechanism leveraging off-the-shelf large multimodal models [17] (see Fig. 1). Specifically, we employ chain-of-thought (CoT)[18] prompting to automatically generate rich, step-by-step natural language weather descriptions for each geographical scene using a single randomly sampled drone-view image. This strategy circumvents manual description and expert controlled, enabling large-scale collection of high-quality and diverse multimodal samples at significantly reduced costs. Moreover, the stepwise reasoning introduced by the CoT prompts ensures semantic accuracy and formatting consistency across generated captions, further enhancing the reliability and usability of descriptions. This ultimately leads to improved generalization for complex and unseen weather conditions. To address the second challenge, we propose a multimodal framework equipped with a text embedding driven dynamic gating mechanism to adaptively reweight and fuse visual features, effectively disentangling scene and weather attributes. Specifically, during training, the framework jointly optimizes multimodal objectives including image-text contrastive (ITC) loss and image-text matching (ITM) loss, aligning drone images with generated weather-aware captions. Additionally, a localized alignment loss is introduced to explicitly enforce consistency between annotated visual regions and textual descriptions across multiple granularities, encouraging the visual encoder to learn robust and weather-invariant scene representations. At inference, visual and textual embeddings are extracted in parallel, and the textual embeddings dynamically modulate visual features via the gating mechanism. The resulting multimodal representation is directly fed into a classification head for localization prediction, avoiding any additional online fine-tuning or dedicated parameter sets for different weather conditions. This design significantly reduces structural complexity and improves deployment efficiency on resource-constrained platforms. The main contributions are as follows:
@@
    State-of-the-Art Generalization: The proposed method has achieved average 87.72% Recall1 on University-1652 and 83.1% on SUS-200 over 10 different weather conditions. For unseen weather combinations (i.e., Dark+Rain+Fog), ours still arrives at 72.15% AP, validating unprecedented cross-domain generalization.
@@ 2 Related Work
Cross-view Geolocalization. Cross-view geo-localization aims to match images captured from different viewpoints with their corresponding geographic locations [1, 19, 20, 21]. Early approaches relied on handcrafted local features such as SIFT [22] and SURF [23], as well as global descriptors like VLAD [24] and Fisher Vector [25], often combined with RANSAC [26] for geometric verification; however, they remain brittle under large viewpoint and illumination changes [27, 28]. With the advent of deep learning, deep metric learning frameworks based on global or part-based contrastive objectives have become dominant [29, 30, 31, 32, 33]. These approaches employ triplet or InfoNCE losses to train end-to-end embeddings and integrate global pooling with spatial attention or multi-region partitioning strategies [34, 35]. Representative examples include the multi-part partitioning scheme of Wang et al. [2], the keypoint attention module of Lin et al. [3], the content-aligned Transformer architecture of Dai et al. [36], self-attention positional encoding by Yang et al. [34], and the dual-path fusion network of Rodrigues et al. [35], all of which substantially enhance cross-view feature alignment. Recent research has begun to address the impact of image degradations such as low-light, motion blur, and synthetic fog, often by using data augmentation, domain adaptation, or cross-modal transformers [37, 38, 39, 40, 41, 42]. However, most existing methods still depend on a limited set of discrete weather labels, restricting generalization to unseen or complex conditions. In contrast, our approach introduces a training-free, all-weather text-guided representation learning framework that leverages open-set weather descriptions to overcome these limitations.
@@
Multi-modality Alignment. In this work, we address weather-aware text-guided representation learning, where the goal is to retrieve drone-view images based on fine-grained weather-related textual cues. Recent advances in vision–language alignment, such as CLIP [43], BLIP [44], and XVLM [16], have established powerful contrastive pre-training and cross-modal attention mechanisms, but prior studies mainly target static semantics or rigid spatial relations [45, 46, 47]. Additional efforts on adaptive fusion and region-word alignment [48, 49, 50, 51] have improved retrieval, yet remain limited by closed vocabularies and overlook dynamic, fine-grained weather semantics. To address these gaps, we propose a framework that generates open-set weather descriptions via Chain-of-Thought prompting and applies text-driven dynamic gating for adaptive feature modulation, achieving robust cross-modal alignment under diverse weather conditions.
@@
Large Vision Language Models for Vision via Prompting. Large Vision Language Models (LVLMs) such as GPT-3/4 [52, 53] and Qwen [17] have recently been applied to vision tasks using prompt engineering. Approaches like VisualGPT [54] and MM-CoT [55] leverage Chain-of-Thought (CoT) prompts [18] to elicit stepwise reasoning for visual question answering and captioning. Yet, scaling LVLMs to cross-view, multi-weather geo-localization remains challenging: free-form text often suffers from hallucinations and lacks semantic or structural consistency [56, 57, 58]. Prior works also overlook prompt design tailored for robust, multi-weather, multi-scale cross-modal alignment. To address this, we introduce the first CoT-driven description pipeline for multi-weather drone-to-satellite geo-localization. Our structured prompts regularize LVLMs’ outputs, enabling scalable generation of high-quality, open-set weather descriptions to advance large-scale vision–language alignment.
@@ 3.1 Open-Weather Description
Single-image Sampling. Since existing benchmarks (e.g., University-1652) are partitioned by geographic regions, we randomly select only one drone-view image per region as a representative to avoid redundancy. Given the prohibitive cost of obtaining real drone-view imagery across diverse weather conditions, we utilize the imgaug [59] library to synthetically realistic weather variations. By parametrically adjusting visibility and occlusion effects such as rain, fog, and snow, we generate high-fidelity, diverse meteorological scenarios that closely resemble real-world conditions.
Weather Estimation Phase. Given a drone-view image with synthetically generated weather, we apply a pretrained large multimodal model [17] for automatic weather description. To mitigate hallucinations, vague phrasing, and inconsistent terminology typical of single-shot captioning, we adopt a Chain-of-Thought prompting strategy inspired by human visual perception. Specifically, we first quantify global visibility, then identify local meteorological cues such as rain-streak reflections or fog diffusion, and finally integrate global and local evidence to determine a reliable weather label. Any description that lacks explicit visibility information, meteorological cues, or contains uncertain terminology (e.g., “possibly” or “uncertain”) is automatically rejected and regenerated. This three-step reasoning pipeline ensures accurate, consistent, and reliable weather descriptions, thus providing a robust prior for the subsequent spatial semantics phase.
@@
Discussion. Existing cross-view geo-localization methods primarily rely on image-based retrieval. While some recent studies [21, 60] integrate textual descriptions, these approaches remain limited to coarse, scene-level labels and largely neglect weather semantics, resulting in compromised robustness under diverse weather conditions. In contrast, we introduce a training-free, CoT-driven pipeline for automatically generating multi-weather semantic descriptions. By employing large multimodal models guided by Chain-of-Thought prompting, we move beyond predefined discrete weather categories and achieve robust generalization to unseen and complex meteorological conditions. Furthermore, our framework systematically annotates both macro-level spatial layouts and micro-level structural attributes, embedding essential spatial cues into the image–text pairs. Relying solely on pretrained multimodal models without manual description, our approach efficiently produces large-scale, semantically consistent multimodal data, significantly facilitating downstream cross-weather multimodal representation learning.
@@ 3.2 Multimodal Alignment Model
which guides the model to hearn subtle semantic distinctions betwwen closely related pairs.
Localized Alignment Module. To enable text-driven fine-grained visual localization, the model responds to both global descriptions $T$ and the region hints $T_{1},T_{2},T_{3}$. For the j-th concept , we extract its [CLS] embedding $x_{\mathrm{cls}}^{j}\in\mathbb{R}^{d}.$ from the cross-encoder, which is projected to normalized coordinate vector via a dedicated two-layer MLP:
@@
where $l^{j}=(c_{x},c_{y},w,h)$ is the ground-truth region, $\hat{l}^{j}$ its prediction, and $\mathrm{IoU}(\cdot,\cdot)$ computes the Intersection-over-Union.
@@ 4 Experiment
Implementation Details. We adopt XVLM [16] as the backbone, whcih is pre-trained on 4M image–caption pairs, integrates BERT [61] as the text encoder and Swin Transformer [62] as the image encoder. The model is optimized using stochastic gradient descent (SGD) [63] with momentum $0.9$ and weight decay $0.0005$. Training 210 epochs, with the learning rate reduced by 0.1 at epoch 120 and by 0.01 at epoch 180. We resize input images to $384\times 384$ pixels and divide them into $32\times 32$ patches. During training, satellite-view images are augmented via random cropping and horizontal flipping, for drone-view images, we first apply style transformations using the imgaug [59] library and then perform the same random crop and flip augmentations. At test time, we compute the Euclidean distance between query and candidate embeddings to measure similarity. All experiments have been implemented in PyTorch [64] and conducted on a single NVIDIA RTX A6000 GPU, with an average inference time of $0.024$s per query.
Dateset. University-1652 [1] is a large-scale cross-view geo-localization dataset comprising images from 1,652 university locations. Each location is represented by satellite, drone, and ground-level images, with 54 drone-view and 1 satellite-view images per building, as well as street-view imagery. The dataset is split into 701 training and 951 test buildings, with no overlap between train and test sets. SUES-200 [65] contains multi-view drone and satellite images from 200 locations in Shanghai, encompassing diverse urban scenes, parks, lakes, and public buildings. Drone images are captured from multiple altitudes (150–300m) to simulate varied real-world conditions.
@@
Zheng et al.[1] [backbone] 67.83 71.74 60.97 65.23 60.29 64.61 55.58 60.09 54.75 59.40 44.85 49.78 57.61 62.03 39.70 44.65 51.85 56.75 58.28 62.83 55.17 59.71
ResNet-101 [66] [backbone] 70.07 73.04 63.87 68.22 63.34 67.59 59.75 64.15 57.45 62.12 48.31 53.28 60.25 64.68 46.12 51.02 56.34 61.23 62.13 66.63 58.76 63.29
DenseNet121 [67] [backbone] 69.48 73.26 64.25 68.47 63.47 67.64 59.29 63.70 59.68 64.13 50.41 55.20 60.21 64.57 48.57 53.41 54.04 58.88 60.74 65.14 59.01 63.44
Swin-T [62] [backbone] 69.27 73.18 66.46 70.52 65.44 69.60 61.79 66.23 63.96 68.21 56.44 61.07 62.68 67.02 50.27 55.18 55.46 60.29 63.81 68.17 61.56 65.95
IBN-Net [68] [backbone] 72.35 75.85 66.68 70.64 67.95 71.73 62.77 66.85 62.64 66.84 51.09 55.79 64.07 68.13 50.72 55.53 57.97 62.52 66.73 70.68 62.30 66.46
LPN [2] [TCSVT’21] 74.33 77.60 69.31 72.95 67.96 71.72 64.90 68.85 64.51 68.52 54.16 58.73 65.38 69.29 53.68 58.10 60.90 65.27 66.46 70.35 64.16 68.14
Sample4Geo* [69] [ICCV’23] 92.70 93.85 88.70 90.55 62.44 66.17 52.76 57.24 52.70 56.77 19.79 23.16 38.19 42.33 46.34 49.91 75.77 78.90 81.54 87.34 61.10 64.32
Safe-Net* [70] [TIP’24] 86.98 88.85 82.12 86.10 67.13 68.90 60.50 63.01 54.80 58.73 32.12 39.77 25.83 26.40 41.10 44.13 69.87 71.15 74.32 76.58 60.48 63.36
CCR* [71] [TCSVT’24] 92.54 93.78 85.57 87.13 67.46 68.82 55.16 59.14 63.11 60.97 27.74 31.48 23.06 46.85 51.10 54.19 75.90 79.16 81.31 87.22 62.30 66.87
MuSe-Net* [29][PR’24] 74.48 77.83 69.47 73.24 70.55 74.14 65.72 69.70 65.59 69.64 54.69 59.24 66.64 70.55 53.85 58.49 61.05 65.51 69.45 73.22 65.15 69.16
Ours 82.78 85.18 81.46 84.03 80.34 83.11 77.60 80.67 78.75 81.69 73.38 76.94 78.41 81.40 67.22 71.06 74.20 77.63 77.26 80.27 77.14 (+11.99+11.99) 80.20 (+11.04+11.04)

### v3

@@ Abstract
Visual geo-localization for drones faces critical degradation under weather perturbations, e.g., rain and fog, where existing methods struggle with two inherent limitations: 1) Heavy reliance on limited weather categories that constrain generalization, and 2) Suboptimal disentanglement of entangled scene-weather features through pseudo weather categories. We present WeatherPrompt, a multi-modality learning paradigm that establishes weather-invariant representations through fusing the image embedding with the text context. Our framework introduces two key contributions: First, a Training-free Weather Reasoning mechanism that employs off-the-shelf large multi-modality models to synthesize multi-weather textual descriptions through human-like reasoning. It improves the scalability to unseen or complex weather, and could reflect different weather strength. Second, to better disentangle the scene and weather features, we propose a multi-modality framework with the dynamic gating mechanism driven by the text embedding to adaptively reweight and fuse visual features across modalities. The framework is further optimized by the cross-modal objectives, including image-text contrastive learning and image-text matching, which maps the same scene with different weather conditions closer in the representation space. Extensive experiments validate that, under diverse weather conditions, our method achieves competitive recall rates compared to state-of-the-art drone geo-localization methods. Notably, it improves Recall@1 by 13.37% under night conditions and by 18.69% under fog and snow conditions. Our code is available at https://github.com/Jahawn-Wen/WeatherPrompt.
@@ 1 Introduction
Drone visual geo-localization aims to match drone-view image with corresponding satellite views, supporting critical applications such as disaster response, urban surveillance, search-and-rescue, and environmental monitoring [Zheng_Wei_Yang_202001, Wang_Zheng_Yan_Zhang_Sun_Zheng_Yang_202202, Lin_Zheng_Zhong_Luo_Li_Yang_Sebe_202203, 968495006, zhu2022demo87, wang2024sparse88]. However, variable weather conditions such as rain, fog and snow introduce noise, occlusions and low visibility, which severely distort image features [feng2024multi04, sindagi2020prior05, liu2022image07] and leading conventional localization methods to suffer drastic performance degradation under extreme weather. Recent advances in cross-modal retrieval show that integrating natural language descriptions can substantially enhance the discrimination power of vision models, allowing better generalization in complex or ambiguous scenarios [yang2023towards08, yan2024dual09, yan2023clip10, vendrow2024inquire11]. Despite this progress, leveraging textual guidance for cross-weather drone geo-localization remains largely underexplored, especially considering the nuanced and dynamic nature of weather conditions encountered in the field. The ability of text to capture complex semantics and fine-grained details [li2022fine64, chow2024unified65] offers a promising avenue for cross-weather generalization. Addressing the all-weather visual geo-localization task presents two primary challenges: (1) Limited Weather Labeling. Existing approaches typically perform domain-specific fine-tuning on a limited set of predefined weather labels (e.g., sunny, rainy). This closed-set paradigm fails to capture the continuous and combinatorial nature of real-world weather, thereby limiting model generalization to unseen or mixed conditions and preventing the exploitation of richer weather semantics. (2) Scene–Weather Feature Entanglement. Existing methods directly inject coarse pseudo-weather labels (e.g., rain, fog) into visual representations during training, leading to a severe entanglement between scene semantics and weather disturbances. Consequently, the model learns suboptimal representations under mixed or unseen weather conditions, fails to disentangle scene content from weather noise, and is severely limited in cross-weather generalization.
@@
Figure 1: Example of the proposed Chain-of-Thought description and matching. Our framework generates structured weather and spatial Text Description via stepwise reasoning. We leverage Off-the-shelf Visual Grounding Model (VGM), i.e., XVLM [xvlm36] to extract local region cues, which are integrated to further refine the matching process. Finally, we match images using weather description, global scene layout, and local region semantics to retrieve the corresponding satellite-view image.
@@
For the first limitation, we propose a training-free weather reasoning mechanism leveraging off-the-shelf large multimodal models [Qwen2.5-VL12] (see Fig. 1). Specifically, we employ chain-of-thought (CoT) [wei2022chain13] prompting to automatically generate rich, step-by-step natural language weather descriptions for each geographical scene using a single randomly sampled drone-view image. This strategy circumvents manual description and expert controlled, enabling large-scale collection of high-quality and diverse multimodal samples at significantly reduced costs. Moreover, the stepwise reasoning introduced by the CoT prompts ensures semantic accuracy and formatting consistency across generated captions, further enhancing the reliability and usability of descriptions. This ultimately leads to improved generalization for complex and unseen weather conditions. To address the second challenge, we propose a multimodal framework equipped with a text embedding-driven dynamic gating mechanism to adaptively reweight and fuse visual features, effectively disentangling scene and weather attributes. Specifically, during training, the framework jointly optimizes multimodal objectives including image-text contrastive (ITC) loss and image-text matching (ITM) loss, aligning drone images with generated weather-aware captions. Additionally, a localized alignment loss is introduced to explicitly enforce consistency between annotated visual regions and textual descriptions across multiple granularities, encouraging the visual encoder to learn robust and weather-invariant scene representations. At inference, visual and textual embeddings are extracted in parallel, and the textual embeddings dynamically modulate visual features via the gating mechanism. The resulting multimodal representation is directly fed into a classification head for localization prediction, avoiding any additional online fine-tuning or dedicated parameter sets for different weather conditions. This design significantly reduces structural complexity and improves the deployment efficiency on resource-constrained platforms. The main contributions are as follows:
@@
    State-of-the-Art Generalization: The proposed method has achieved average 87.72% Recall@1 on University-1652 and 83.1% on SUS-200 over 10 different weather conditions. For unseen weather combinations (i.e., Dark+Rain+Fog), ours still arrives at 72.15% AP, validating unprecedented cross-domain generalization.
@@ 2 Related Work
Cross-view Geolocalization. Cross-view geo-localization aims to match images captured from different viewpoints with their corresponding geographic locations [Zheng_Wei_Yang_202001, berton2022deep14, jin2017learned15, chu2024towards16]. Early approaches relied on hand-crafted local features such as SIFT [cruz2012scale17] and SURF [bay2006surf18], as well as global descriptors like VLAD [zhang2021vector19] and Fisher Vector [sanchez2013image20], often combined with RANSAC [barath2022learning21] for geometric verification; however, they remain brittle under large viewpoint and illumination changes [Castaldo_Zamir_Angst_Palmieri_Savarese_201522, Lin_Belongie_Hays_201323]. With the advent of deep learning, deep metric learning frameworks based on global or part-based contrastive objectives have become dominant [wang2024Muse24, Workman_Souvenir_Jacobs_201525, Lin_YinCui_Belongie_Hays_201526, vepa2024integrating66, choi2023depth68, zheng2017discriminatively]. These approaches employ triplet or InfoNCE losses to train end-to-end embeddings and integrate global pooling with spatial attention or multi-region partitioning strategies [yang2021cross27, rodrigues2022global28]. Representative examples include the multi-part partitioning scheme et al. [Wang_Zheng_Yan_Zhang_Sun_Zheng_Yang_202202], the keypoint attention module [Lin_Zheng_Zhong_Luo_Li_Yang_Sebe_202203], the content-aligned Transformer architecture [dai2021transformer29], self-attention positional encoding by Yang et al. [yang2021cross27], the dual-path fusion network [rodrigues2022global28], and Bird’s Eye View (BEV) [ju2024video2bev], all of which substantially enhance cross-view feature alignment. Recent research has begun to address the impact of image degradations such as low-light, motion blur, and synthetic fog, often by using data augmentation, domain adaptation, or cross-modal transformers [tan2023mapd31, ma2022both30, zhang2023cross32, zhu2022transgeo33, qu2024lush69, chen2024restoreagent70]. However, most existing methods still depend on a limited set of discrete weather labels, restricting generalization to unseen or complex conditions. In contrast, our approach introduces a training-free, all-weather text-guided representation learning framework that leverages open-set weather descriptions to overcome these limitations.
@@
Multi-modality Alignment. In this work, we address weather-aware text-guided representation learning, where the goal is to retrieve drone-view images based on fine-grained weather-related textual cues. Recent advances in vision–language alignment, such as CLIP [radford2021learning34], BLIP [li2022blip35], and XVLM [xvlm36], have established powerful contrastive pre-training and cross-modal attention mechanisms, but previous studies mainly target static semantics or rigid spatial relations [zhu2024mvp37, chen2024rsprompter38, yuan2023parameter39]. Additional efforts on adaptive fusion and region-word alignment [zheng2020dual40, wang2019camp41, chen2020uniter43, li2020oscar44] have improved retrieval, but remain limited by closed vocabularies and overlook dynamic, fine-grained weather semantics. To address these gaps, we propose a framework that generates open-set weather descriptions via Chain-of-Thought prompting and applies text-driven dynamic gating for adaptive feature modulation, achieving robust cross-modal alignment under diverse weather conditions.
@@
Large Vision Language Models for Vision via Prompting. Large Vision Language Models (LVLMs) such as GPT‑3/4 [brown2020language45, achiam2023gpt46] and Qwen [Qwen2.5-VL12] have recently been applied to vision tasks using prompt engineering. Approaches like VisualGPT [chen2022visualgpt47] and MM‑CoT [zhang2023multimodal48] leverage Chain-of-Thought (CoT) prompts [wei2022chain13] to elicit stepwise reasoning for visual question answering and captioning. However, scaling LVLMs to cross-view, multi-weather geo-localization remains challenging: free-form text often suffers from hallucinations and lacks semantic or structural consistency [sriramanan2024llm49, jiang2024hallucination50, kim2024exploiting51]. Prior works also overlook prompt design tailored for robust, multi-weather, multi-scale cross-modal alignment. To address this, we introduce the first CoT-driven description pipeline for multi-weather drone-to-satellite geo-localization. Our structured prompts regularize LVLMs’ outputs, enabling scalable generation of high-quality, open-set weather descriptions to advance large-scale vision–language alignment.
@@ 3.1 Open-Weather Description
Single-image Sampling. Since existing benchmarks (e.g., University-1652) are partitioned by geographic regions, we randomly select only one drone-view image per region as a representative to avoid redundancy. Given the prohibitive cost of obtaining real drone-view imagery across diverse weather conditions, we utilize the imgaug [imgaug67] library to synthesize realistic weather variations. By parametrically adjusting visibility and occlusion effects such as rain, fog, and snow, we generate high-fidelity, diverse meteorological scenarios that closely resemble real-world conditions.
Weather Estimation Phase. Given a drone-view image with synthetically generated weather, we apply a pretrained large multimodal model [Qwen2.5-VL12] for automatic weather description. To mitigate hallucinations, vague phrasing, and inconsistent terminology typical of single-shot captioning, we adopt a Chain-of-Thought prompting strategy inspired by human visual perception. Specifically, we first quantify global visibility, then identify local meteorological cues such as rain-streak reflections or fog diffusion, and finally integrate global and local evidence to determine a reliable weather label. Any description that lacks explicit visibility information, meteorological cues, or contains uncertain terminology (e.g., “possibly” or “uncertain”) is automatically rejected and regenerated. This three-step reasoning pipeline ensures accurate, consistent, and reliable weather descriptions, thus providing a robust prior for the subsequent spatial semantics phase.
@@
Discussion. Existing cross-view geo-localization methods primarily rely on image-based retrieval. While some recent studies [chu2024towards16, ye2024cross52] integrate textual descriptions, these approaches remain limited to coarse, scene-level labels and largely neglect weather semantics, resulting in compromised robustness under diverse weather conditions. In contrast, we introduce a training-free, CoT-driven pipeline for automatically generating multi-weather semantic descriptions. By employing large multimodal models guided by Chain-of-Thought prompting, we move beyond predefined discrete weather categories and achieve robust generalization to unseen and complex meteorological conditions. Furthermore, our framework systematically annotates both macro-level spatial layouts and micro-level structural attributes, embedding essential spatial cues into the image–text pairs. Relying solely on pretrained multimodal models without manual description, our approach efficiently produces large-scale, semantically consistent multimodal data, significantly facilitating downstream cross-weather multimodal representation learning.
@@ 3.2 Multimodal Alignment Model
which guides the model to hear subtle semantic distinctions between closely related pairs.
Localized Alignment Module. To enable text-driven fine-grained visual localization, the model responds to both global descriptions $T$ and the region hints $T_{1},T_{2},T_{3}$. For the j-th concept , we extract its [CLS] embedding $x_{\mathrm{cls}}^{j}\in\mathbb{R}^{d}.$ from the cross-encoder, which is projected to normalized coordinate vector via a dedicated two‐layer MLP:
@@
where $l^{j}=(c_{x},c_{y},w,h)$ is the ground‐truth region, $\hat{l}^{j}$ its prediction, and $\mathrm{IoU}(\cdot,\cdot)$ computes the Intersection‐over‐Union.
@@ 4 Experiment
Implementation Details. We adopt XVLM [xvlm36] as the backbone, which is pre-trained on 4M image–caption pairs, integrates BERT [devlin2019bert54] as the text encoder and Swin Transformer [liu2021swin55] as the image encoder. The model is optimized using stochastic gradient descent (SGD) [robbins1951stochastic53] with momentum $0.9$ and weight decay $0.0005$. Training 210 epochs, with the learning rate reduced by 0.1 at epoch 120 and by 0.01 at epoch 180. We resize input images to $384\times 384$ pixels and divide them into $32\times 32$ patches. During training, satellite‐view images are augmented via random cropping and horizontal flipping, for drone‐view images, we first apply style transformations using the imgaug [imgaug67] library and then perform the same random crop and flip augmentations. At test time, we compute the Euclidean distance between query and candidate embeddings to measure similarity. All experiments have been implemented in PyTorch [paszke2019pytorch56] and conducted on a single NVIDIA RTX A6000 GPU, with an average inference time of $0.024$s per query.
Dateset. University-1652 [Zheng_Wei_Yang_202001] is a large-scale cross-view geo-localization dataset comprising images from 1,652 university locations. Each location is represented by satellite, drone, and ground-level images, with 54 drone-view and 1 satellite-view images per building, as well as street-view imagery. The dataset is split into 701 training and 951 test buildings, with no overlap between train and test sets. SUES-200 [zhu2023sues57] contains multi-view drone and satellite images from 200 locations in Shanghai, encompassing diverse urban scenes, parks, lakes, and public buildings. Drone images are captured from multiple altitudes (150–300m) to simulate varied real-world conditions.
@@
Zheng et al.[Zheng_Wei_Yang_202001] [backbone] 67.83 71.74 60.97 65.23 60.29 64.61 55.58 60.09 54.75 59.40 44.85 49.78 57.61 62.03 39.70 44.65 51.85 56.75 58.28 62.83 55.17 59.71
ResNet-101 [He_Zhang_Ren_Sun_201658] [backbone] 70.07 73.04 63.87 68.22 63.34 67.59 59.75 64.15 57.45 62.12 48.31 53.28 60.25 64.68 46.12 51.02 56.34 61.23 62.13 66.63 58.76 63.29
DenseNet121 [huang2017densely59] [backbone] 69.48 73.26 64.25 68.47 63.47 67.64 59.29 63.70 59.68 64.13 50.41 55.20 60.21 64.57 48.57 53.41 54.04 58.88 60.74 65.14 59.01 63.44
Swin-T [liu2021swin55] [backbone] 69.27 73.18 66.46 70.52 65.44 69.60 61.79 66.23 63.96 68.21 56.44 61.07 62.68 67.02 50.27 55.18 55.46 60.29 63.81 68.17 61.56 65.95
IBN-Net [Pan_Luo_Shi_Tang_201860] [backbone] 72.35 75.85 66.68 70.64 67.95 71.73 62.77 66.85 62.64 66.84 51.09 55.79 64.07 68.13 50.72 55.53 57.97 62.52 66.73 70.68 62.30 66.46
LPN [Wang_Zheng_Yan_Zhang_Sun_Zheng_Yang_202202] [TCSVT’21] 74.33 77.60 69.31 72.95 67.96 71.72 64.90 68.85 64.51 68.52 54.16 58.73 65.38 69.29 53.68 58.10 60.90 65.27 66.46 70.35 64.16 68.14
Sample4Geo* [deuser2023sample4geo61] [ICCV’23] 92.70 93.85 88.70 90.55 62.44 66.17 52.76 57.24 52.70 56.77 19.79 23.16 38.19 42.33 46.34 49.91 75.77 78.90 81.54 87.34 61.10 64.32
Safe-Net* [lin2024self62] [TIP’24] 86.98 88.85 82.12 86.10 67.13 68.90 60.50 63.01 54.80 58.73 32.12 39.77 25.83 26.40 41.10 44.13 69.87 71.15 74.32 76.58 60.48 63.36
CCR* [du2024ccr63] [TCSVT’24] 92.54 93.78 85.57 87.13 67.46 68.82 55.16 59.14 63.11 60.97 27.74 31.48 23.06 46.85 51.10 54.19 75.90 79.16 81.31 87.22 62.30 66.87
MuSe-Net* [wang2024Muse24][PR’24] 74.48 77.83 69.47 73.24 70.55 74.14 65.72 69.70 65.59 69.64 54.69 59.24 66.64 70.55 53.85 58.49 61.05 65.51 69.45 73.22 65.15 69.16
[…]